Circular No. 5754

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 T. Balonek, Colgate University, communicates: “SN 1993J appears to

have reached a minimum on Apr. 5 UT, and is now once again increasing in brightness. Preliminary reductions of CCD images taken with the Burrell Schmidt telescope on Kitt Peak indicate that the supernova is 0.11 mag brighter in V on Apr. 7.15 than on Apr. 6.15, 0.09 mag brighter in B, 0.01 brighter in U, 0.10 brighter in R, and 0.08 brighter in I (errors about ± 0.02-0.03).”

G. Sonneborn, Goddard Space Flight Center; P. M. Rodriguez, W. Wamsteker, IUE Observatory, European Space Agency; C. Fransson, Stockholm Observatory; and R. P. Kirshner, Center for Astrophysics, report: “Ultraviolet spectra (range 115-320 nm, resolution 0.5 nm) of SN 1993J have been obtained daily through Apr. 5.7 UT with the IUE satellite. The ultraviolet continuum flux at 127.5, 173.5, 245.0, and 290.0 nm (5.0-nm bandpasses) has decreased by factors of 203, 83, 29, and 11, respectively, between Mar. 30.2 and Apr. 4.6 (cf. IAUC 5738). The time of maximum light appears to have been prior to Mar. 30.2 at all ultraviolet wavelengths. The ultraviolet flux changes were initially rapid, but have slowed considerably since Apr. 2. During this period, the spectrum has remained devoid of significant spectral features, aside from interstellar absorption lines and N V 124.0-nm emission. Such a rapid decrease is not unprecedented. For example, in SN 1987A the ultraviolet continuum (< 160 nm) faded by more than a factor of 1000 between 1987 Feb. 24 and 28. The N V emission line flux measured in the low-resolution spectra has followed a light curve similar to the ultraviolet continuum, decreasing by a factor of 43 during this same time interval; the integrated N V flux was 3.9 × 10^{−12} erg cm^{−2} s^{−1} on Mar. 30.2. The properties of the N V line (narrow width, large flux, prompt appearance, rapid decay) suggest that this emission arises in slowly moving, dense material close to the progenitor, most likely a stellar wind, characteristic of late-type supergiants. The high state of ionization strongly indicates that the wind has been ionized and heated by the ultraviolet and x-rays from the shock breakout, as well as the radiation resulting from the interaction of the ejecta and circumstellar medium (e.g., Lundqvist and Fransson 1992, A.Ap. 192, 221). The interaction of the supernova ejecta with stellar wind material will create one high-velocity shock and one reverse shock, producing soft x-ray and radio emission, consistent with recently reported detections (IAUC 5750, 5752).”

1993 April 7 Daniel W. E. Green